— — Multinomická věta

Multinomická věta

∀n ∈ ℕ_{0}, ∀m ∈ ℕ, ∀x_{1}, …, x_{m} ∈ ℝ :

(x_{1} + x_{2} + ⋯ + x_{m})^{n} = ∑_{{k_{1} ,k_{2} ,...,k_{m}≥0 \atop k_{1} +k_{2} +···+k_{m}=n}} \binom{n}{k_{1}, k_{2}, …, k_{m}}(x_{1}^{k_{1}}x_{2}^{k_{2}} ⋯ x_{m}^{k_{m}})

Bez důkazu.

Poznámka:

Pro m = 2 odpovídá binomické větě.

Multinomický koeficient

\binom{n}{k_{1}, …, k_{m}} = \frac{n!}{k_{1}!k_{2}! ⋯ k_{n}!}

Počet způsobů zařazení čísel 1, …, n do m množin x_{1}, …, x_{m} tak, aby |x_{1}| = k_{1}, …, |x_{m}| = k_{m}.

Platí

\binom{n}{k} = \binom{n}{k, n − k}

Příklad:

200 dětí, 3 autobusy (80, 70, 50), počet možností rozmístění

\binom{200}{80, 70, 50}

/L:1 1